suppose the amplitude of a simple pendulum having a bob of mass m.then find the tension in string at its mean position.

Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

let 'α' radians be the amplitude. at mean position, kinetic energy=a(1-cos(alpha)),a is length of pendulum

tension in string at mean position=

mg+m(v^2)/a   m(v^2)/2=a(1-cos(α))